# FIT-GNN: Faster Inference Time for GNNs that ‘FIT’ in Memory Using Coarsening

### v4

@@ 5 Experiments
  • Gc-train-to-Gs-train: Train the GNN model on $G^{\prime}$ as per Algorithm 3, then use the learned weight as an initialization for subgraph-level training and final inference on $\mathcal{G}_{s}$. Our aim was to examine whether pretraining on $G^{\prime}$ (which requires less time) followed by fine-tuning on $\mathcal{G}_{s}$ for fewer epochs can improve performance.
@@
Dataset descriptions and details related to hyperparameters, device configuration, and key packages used for the experiments are in Section D and E. Source code: FIT-GNN.
Table 2: Datasets and corresponding splits used in our experiments. ‘c’ denotes the number of classes in node classification datasets. Public split “fixed” from 10.5555/3045390.3045396 is used for Cora, Citeseer, and PubMed. For the “random” split, we use 20 nodes per class for training, 30 per class for validation, and the rest for testing. For QM9, we follow DBLP:journals/corr/GilmerSRVD17 and predict one property from each broad category: dipole moment ($\mu$), HOMO-LUMO gap ($\epsilon_{\mathrm{HOMO}}-\epsilon_{\mathrm{LUMO}}$), zero-point vibration energy (ZPVE), and atomization energy ($U^{\mathrm{ATOM}}$) at $298.15\mathrm{K}$.
@@
Node Regression Chameleon rozemberczki2019multiscale 30% 20% 50%
Crocodile rozemberczki2019multiscale 30% 20% 50%
Squirrel rozemberczki2019multiscale 30% 20% 50%
Node Classification Cora McCallum2000 $20\times c$ $30\times c$ Rest
Citeseer 10.1145/276675.276685 $20\times c$ $30\times c$ Rest
Pubmed Sen_Namata_Bilgic_Getoor_Galligher_Eliassi-Rad_2008 $20\times c$ $30\times c$ Rest
Coauthor Physics DBLP:journals/corr/abs-1811-05868 $20\times c$ $30\times c$ Rest
DBLP 10.1145/1401890.1402008 $20\times c$ $30\times c$ Rest
OGBN-Products 10.5555/3495724.3497579 $20\times c$ $30\times c$ Rest
Graph Regression QM9 DBLP:journals/corr/WuRFGGPLP17 50% 25% 25%
ZINC (Subset) DBLP:journals/corr/Gomez-Bombarelli16 50% 25% 25%
Graph Classification PROTEINS Morris+2020 50% 25% 25%
AIDS Morris+2020 50% 25% 25%
@@ 6.1 Model Performance
Chameleon Crocodile Squirel
@@
Methods Models
Reduction
Ratio (r)
Dataset
Cora Citeseer Pubmed DBLP Physics
0.5 $0.712\pm 0.059$ $0.501\pm 0.085$ $0.409\pm 0.044$ $0.464\pm 0.030$ OOM
Node Regression: Table 3 shows the node regression error (mean and standard deviation of the lowest $10$ MAE). As the coarsening ratio increases, more subgraphs are created, and the size of each subgraph also reduces, and important neighboring node information gets lost. Hence, an increase in the coarsening ratio decreases performance.
Full $0.873\pm 0.000$
SGGC OOM
GCOND OOM
BONSAI OOM
FIT-GNN $\mathbf{0.894\pm 0.000}$
Node Classification: Table 4 shows the results for node classification accuracy (mean and standard deviation of the top $10$ accuracies) with coarsening ratios $0.3$ and $0.5$ using the Cluster Nodes approach. Comparison with other coarsening ratios is in Table 10. Results show that the accuracy is comparable to the baselines. Figure 5 shows the comparison of performance between different experimental setups for different methods of appending nodes on the Cora dataset. From the Figure, we observe that Gc-train-to-Gs-train offers the best result among the three experimental setups for the node classification task. For the ’None’ method of appending nodes, where no additional nodes are added to each subgraph, we notice that it performs poorly compared to other methods, indicating that the Extra Node and Cluster Node methods indeed provide additional information for better empirical performance. We also observe in the ’None’ method that, with an increase in the coarsening ratio, the performance drops. This is because, as the coarsening ratio increases, the size of each subgraph gets smaller, hence more edges are cut due to partitioning, leading to more information loss. For the Gc-train-to-Gs-infer method, we observe that the performance drops as the coarsening ratio increases. This can be argued because of the change in the distribution of the coarsening graph and the set of subgraphs. When training on $G_{s}$, the distribution in the train and test sets is similar, hence attaining better performance. While comparing the methods of appending nodes, we infer that the Cluster Node method is performing similarly compared to the Extra Nodes method, if not better. Table 5 shows results on the OGBN-Products dataset. Result on training on the full dataset from the OGBN-Products leaderboard [Link]. For training on the full dataset, $8\%$ for training, next top $2\%$ for validation, and the rest for testing. For our approach, we use $0.038\%$ training, $0.057\%$ validation, and the rest for testing.
@@
ZINC (Subset) QM9 (prop 0) QM9 (prop 4) QM9 (prop 6) QM9 (prop 13)
@@
Graph Regression: Table 6 shows the results for the graph regression task on ZINC (Subset) and QM9 dataset. All the results show that the FIT-GNN model’s performance is better than the baselines. It is also observed that a lower coarsening ratio yields better loss, which implies that the model performs better on molecular graphs when the subgraph size is large. It implies that finer subgraphs lose out on global information about the molecule, which is necessary for the prediction. Similarly, when inferred on a coarsened graph for each graph in the dataset, the performance degrades. This is because, when a graph is reduced to a smaller graph, the positional information is lost, which is crucial since in graphs representing different molecules, the position of a certain component in a molecule results in different properties of the molecule.
@@
Methods Models
Reduction
Ratio (r)
Dataset
@@
Graph Classification: Table 7 shows graph classification results on AIDS and PROTEINS datasets, where training and inference are done on $G^{\prime}$. For baselines like DOSCOND and KIDD, we use ‘Graph per class’, whereas we use ‘Reduction Ratio (r)’ as a metric to denote how these algorithms reduce the size of training data. DOSCOND and KIDD propose to generate synthetic graphs that mimic the training data distribution. Therefore, to cover all the graphs in the dataset appropriately, this approach generates {1, 10, 50} graphs per class. However, the graph size is kept similar to the graph size in the training set. These approaches also don’t cover the test set. Compared to these baselines, our approach reduces each graph into a set of subgraphs and a coarsened graph based on a reduction ratio. As a result, each graph in the training set and test set gets reduced. From our results, we see that FIT-GNN outperforms all the baselines.
Table 11 and 12 present the ablation study on various coarsening algorithms for all the aforementioned tasks. From the results, we observe that variation_neighborhoods coarsening algorithm yields consistently better results compared to others. Therefore, we report results on all datasets using variation_neighborhoods coarsening algorithm.
@@ 6.2 Inference Time and Memory
Table 8: Inference time (sec) (Lower is better) comparison for Node-Level tasks and Graph-Level tasks. We show single-node prediction time using two different coarsening ratios for Node-Level tasks. Here, we used Cluster Nodes for the node-level task. We used Gc-train-to-Gc-infer experimental setup for the graph-level tasks.
@@
(a) Node-Level Tasks
@@
Dastaset Baselines $\mathbf{r=0.3}$ $\mathbf{r=0.5}$
@@
In Table 8(a), we show the average time to predict for $1000$ queries (Prediction for Node $v_{i}$) in the Baseline and FIT-GNN models for node regression and classification tasks. The baseline model processes the entire graph, increasing inference time, especially for large graphs. In contrast, the FIT-GNN model only requires the relevant subgraph, resulting in faster predictions. For larger datasets like the subset of OGBN-Products, baseline inference is not feasible with our computational capacity. Therefore, to show a comparison, we take a subset of OGBN-Products with $165000$ nodes and $4340428$ edges. We see up to $\mathbf{100\times}$. Section C elaborates more on the inference time for datasets with a larger number of nodes. Traditional methods fail to infer for very large datasets due to memory constraints. However, our approach can give an inference on the whole graph. Figure 3 and Table 13 show the detailed comparison of memory consumption for different datasets for both Extra Nodes and Cluster Nodes methods, along with the baseline memory, which highlights that the FIT-GNN model uses up to $\mathbf{100\times}$ less memory than the Baseline. Figure 6 shows the time taken to coarsen for different coarsening ratios for the Cora dataset. As the coarsening ratio increases, the time also increases, since the number of subgraphs increases with the coarsening ratio.
@@ 7 Conclusion
In this paper, we have focused on inference time and memory and presented a new way to utilize existing graph coarsening algorithms for GNNs. We have provided theoretical insight corresponding to the number of nodes in the graph for which the FIT-GNN model reduces the time and space complexity. Empirically, we have shown that our method is comparable to the uncoarsened Baseline while being orders of magnitude faster in terms of inference time and consuming a fraction of the memory. A few possible future directions involve studying directed and weighted graphs, focusing on the theoretical connections between Extra Nodes and Cluster Nodes, and tackling real-world data-intensive challenges, for example, weather forecasting.
@@ A.1 Extra Nodes
  • Nodes with 1-hop and 2-hop neighbours in $G_{i}$. Let us call this set $S_{1}$
  • Nodes with 1-hop neighbours in $G_{i}$ but $\exists v$ in 2-hop neighbourhood that is not in $G_{i}$. Let us call this set $S_{2}$
  • Node where $\exists v$ in 1-hop and 2-hop neighbourhood that is not in $G_{i}$. Let us call this set $S_{3}$.
@@ A.2 Cluster Nodes
Given a partition matrix $P$, the features of the coarsened node are $X^{\prime}=P^{T}X$. Given a normalized partition matrix, the features of a node $v^{\prime}_{i}\in G^{\prime}$ is the degree-weighted average of the features of nodes in $C_{i}$. This is one of the functions $f$ used to create the features of the cluster node from $C_{i}$.
@@ Appendix B More Algorithms
Algorithm 3 Training GNN (Train on $G^{\prime}$) (10.1145/3447548.3467256)
@@
4:Construct feature matrix for $G^{\prime}$ by $X^{\prime}=P^{T}X$;
5:Construct labels for $G^{\prime}$ by $Y^{\prime}=\arg\max(P^{T}Y)$ ;
@@
3:  $X^{(i)}=\sigma(\tilde{D}^{-\frac{1}{2}}\tilde{A}\tilde{D}^{-\frac{1}{2}}X^{(i-1)}\mathcal{W}^{(i-1)})$ $\quad\#$ Equation 1
@@
[…]
@@ Appendix C More on Time and Space complexity
Figure 4: The figure shows the feasibility of our inference methods for different node-level datasets against coarsening ratios. Two important inference setups for each dataset: single-node inference with the computational cost $\mathcal{O}\left(\max_{i}(\bar{n_{i}}^{2}d+\bar{n_{i}}d^{2})+n\right)$, and full-graph inference with computational cost $\mathcal{O}\left(\sum_{i=1}^{k}(\bar{n_{i}}^{2}d+\bar{n_{i}}d^{2})\right)$, are compared against the baseline (classical GNN) computational cost of $\mathcal{O}\left(n^{2}d+nd^{2}\right)$. The coarsening method used is variation_neighborhoods. Both axes are in logarithmic scale (base 10).
@@
    We observe lower inference time complexity of Cluster_node method as compared to Extra_node method under both inference setups. This is expected as $|C_{G_{i}}|\leq|\mathcal{E}_{{G_{i}}}|,\ \forall i\in\{1,2,\dots,k\}$.

### v5

@@ 5 Experiments
  • Gc-train-to-Gs-train: Train the GNN model on $G^{\prime}$ as per Algorithm 3, then use the learned weight as an initialization for subgraph-level training and final inference on $\mathcal{G}_{s}$. We examine whether pretraining on $G^{\prime}$ (which requires less time) followed by fine-tuning on $\mathcal{G}_{s}$ for fewer epochs can improve performance.
@@
Dataset descriptions and details related to hyperparameters, device configuration, and key packages used for the experiments are in Section D and E. Source code: https://github.com/Roy-Shubhajit/FIT-GNN.
Table 2: Datasets and corresponding splits used in our experiments. ‘c’ denotes the number of classes in node classification datasets. Public split “fixed” by Yang et al. (2016) is used for Cora, Citeseer, and PubMed. For the “random” split, we use 20 nodes per class for training, 30 per class for validation, and the rest for testing. For QM9, we follow Gilmer et al. (2017) and predict one property from each broad category: dipole moment ($\mu$), HOMO-LUMO gap ($\epsilon_{\mathrm{HOMO}}-\epsilon_{\mathrm{LUMO}}$), zero-point vibration energy (ZPVE), and atomization energy ($U^{\mathrm{ATOM}}$) at $298.15\mathrm{K}$.
@@
Node Regression Chameleon (Rozemberczki et al., 2021) 30% 20% 50%
Crocodile (Rozemberczki et al., 2021) 30% 20% 50%
Squirrel (Rozemberczki et al., 2021) 30% 20% 50%
Node Classification Cora (McCallum et al., 2000) $20\times c$ $30\times c$ Rest
Citeseer (Giles et al., 1998) $20\times c$ $30\times c$ Rest
Pubmed (Sen et al., 2008) $20\times c$ $30\times c$ Rest
Coauthor Physics (Shchur et al., 2018) $20\times c$ $30\times c$ Rest
DBLP (Tang et al., 2008) $20\times c$ $30\times c$ Rest
OGBN-Products (Hu et al., 2020) $20\times c$ $30\times c$ Rest
Graph Regression QM9 (Wu et al., 2018) 50% 25% 25%
ZINC (Subset) (Gómez-Bombarelli et al., 2018) 50% 25% 25%
Graph Classification PROTEINS (Morris et al., 2020) 50% 25% 25%
AIDS (Morris et al., 2020) 50% 25% 25%
@@ 6.1 Model Performance
Full (Luo et al., 2024) $0.823\pm 0.001$
SGGC OOM
GCOND OOM
BONSAI OOM
FIT-GNN $\mathbf{0.894\pm 0.000}$
Methods Models Reduction Ratio (r) Dataset
Cora Citeseer Pubmed DBLP Physics
0.5 $0.712\pm 0.059$ $0.501\pm 0.085$ $0.409\pm 0.044$ $0.464\pm 0.030$ OOM
@@
Chameleon Crocodile Squirrel
@@
Node Classification: Table 4 shows the results for node classification accuracy (mean and standard deviation of the top $10$ accuracies) with coarsening ratios $0.3$ and $0.5$ using the Cluster Nodes approach. Comparison with other coarsening ratios is in Table 12. Results show that the accuracy is comparable to the baselines. Figure 3 shows the comparison of performance between different experimental setups for different methods of appending nodes on the Cora dataset. From the Figure, we observe that Gc-train-to-Gs-train offers the best result among the three experimental setups for the node classification task. For the ‘None’ method of appending nodes, where no additional nodes are added to each subgraph, we notice that it performs poorly compared to other methods, indicating that the Extra Node and Cluster Node methods indeed provide additional information for better empirical performance. We also observe in the ‘None’ method that, with an increase in the coarsening ratio, the performance drops. This is because, as the coarsening ratio increases, the size of each subgraph gets smaller, hence more edges are cut due to partitioning, leading to more information loss. For the Gc-train-to-Gs-infer method, we observe that the performance drops as the coarsening ratio increases. This can be argued because of the change in the distribution of the coarsened graph and the set of subgraphs. When training on $G_{s}$, the distribution in the train and test sets is similar, hence attaining better performance. While comparing the methods of appending nodes, we infer that the Cluster Node method is performing similarly compared to the Extra Nodes method, if not better. Table 3 shows results on the OGBN-Products dataset. We use the result by Luo et al. (2024) for the full graph setting, which used the OGBN’s standard split, while the split used in our work is detailed in Table 2.
Node Regression: Table 5 presents the node regression error (mean and standard deviation of the lowest 10 MAE). Interestingly, our results demonstrate that utilizing localized subgraphs for inference (FIT-GNN) significantly improves performance compared to full-graph inference. To investigate this counterintuitive performance leap, we conducted a rigorous ablation study (Appendix G). Our analysis indicates that this improvement is driven by the structural properties of the input data during inference rather than the training regime itself. First, partitioning the graph creates localized contexts that are statistically more homogeneous; we observed that the label standard deviation within individual subgraphs is drastically lower than the global variation, presenting a simpler optimization landscape for the model. Second, while coarsening (e.g., at a ratio of $r=0.5$) causes a vast majority of nodes to lose a significant portion of their distant 2nd-hop neighborhood, this structural loss actually acts as implicit adversarial pruning. In these specific heterophilic graphs, long-range information often introduces noise or adversarial signals. By filtering out this noise through coarsening, the model is able to fully exploit the low-variance local structures, ultimately leading to the observed reduction in regression error.
@@
ZINC (Subset) QM9 ($\mu$) QM9 ($\Delta\epsilon$) QM9 (ZPVE) QM9 ($U^{\textup{ATOM}}$)
@@
Graph Regression: Table 6 shows the results for the graph regression task on ZINC (Subset) and QM9 dataset. All the results show that the FIT-GNN model’s performance is better than the baselines. It is also observed that a lower coarsening ratio yields better loss, which implies that the model performs better on molecular graphs when the subgraph size is large. We observe that finer subgraphs lose out on global information about the molecule, which is necessary for the prediction. Similarly, the performance degrades in the Gc-train-to-Gc-infer setup, where we infer on the coarsened graphs. This is because, when a graph is reduced to a smaller graph, the individual node information is lost, which is crucial since, in graphs representing different molecules, components formed by different atoms result in different properties of the molecule.
Graph Classification: Table 7 shows graph classification results on AIDS and PROTEINS datasets, where training and inference are done on $G^{\prime}$. For baselines like DOSCOND and KIDD, we use ‘Graph per class’, whereas we use ‘Reduction Ratio (r)’ as a metric to denote how these algorithms reduce the size of training data. DOSCOND and KIDD propose to generate synthetic graphs that mimic the training data distribution. Therefore, to cover all the graphs in the dataset appropriately, this approach generates {1, 10, 50} graphs per class. However, the graph size is kept similar to the graph size in the training set. These approaches also don’t cover the test set. Compared to these baselines, our approach transforms each graph into a set of subgraphs and a coarsened graph based on a reduction ratio. As a result, each graph in the training set and test set gets reduced. From our results, we see that FIT-GNN outperforms all the baselines.
Table 14 and 15 present the ablation study on various coarsening algorithms for all the aforementioned tasks. From the results, we observe that variation_neighborhoods coarsening algorithm yields consistently better results compared to others.
@@
Methods Models Reduction Ratio (r) Dataset
@@ 6.2 Inference Time and Memory
Table 8: Inference time (sec) (Lower is better) comparison for Node-Level tasks and Graph-Level tasks. We show single-node prediction time using two different coarsening ratios for Node-Level tasks. Here, we used Cluster Nodes. We used Gc-train-to-Gc-infer experimental setup for the graph-level tasks.
(a) Node-Level Tasks
@@
(b) Graph-Level Tasks
@@
Dataset Baselines $\mathbf{r=0.3}$ $\mathbf{r=0.5}$
@@
In Table 8(a), we show the average time to predict for $1000$ queries (Prediction for Node $v_{i}$) in the Baseline and FIT-GNN models for node regression and classification tasks. The baseline model processes the entire graph, increasing inference time, especially for large graphs. In contrast, the FIT-GNN model only requires the relevant subgraph, resulting in faster predictions. For larger datasets like OGBN-Products, baseline inference is not feasible with our computational capacity. Therefore, to show a comparison, we take a subset of OGBN-Products with $165000$ nodes and $4340428$ edges with respect to which we see up to $\mathbf{100\times}$ speedup in inference time. Section C elaborates more on the inference time for datasets with a larger number of nodes. Traditional methods fail to infer for very large datasets due to memory constraints. However, our approach enables inference on the whole graph. Figure 4 and Table 13 show the detailed comparison of memory consumption for different datasets for both Extra Nodes and Cluster Nodes methods, along with the baseline memory, which highlights that the FIT-GNN model uses up to $\mathbf{100\times}$ less memory than the Baseline. Figure 6 shows the time taken to coarsen for different coarsening ratios for the Cora dataset. As the coarsening ratio increases, the time also increases, since the number of subgraphs increases.
@@ 7 Conclusion
In this paper, we have focused on inference time and memory and presented a new way to utilize existing graph coarsening algorithms for GNNs. We have provided theoretical insight corresponding to the number of nodes in the graph for which the FIT-GNN model reduces the time and space complexity. Empirically, we have shown that our method is comparable to the uncoarsened Baseline while being orders of magnitude faster in terms of inference time and consuming a fraction of the memory. Future directions include extending the FIT-GNN methodology to directed, weighted, heterogeneous, and temporal graphs, alongside tackling data-intensive real-world challenges like weather forecasting. Theoretically, we aim to quantify the information loss for the Cluster Node method, investigate its connection to Extra Nodes, and explore novel strategies beyond both to further mitigate graph partitioning loss.
8 Acknowledgment
Anirban Dasgupta acknowledges the support by SERB MATRICS and SERB CRG grants and the support from N Rama Rao Chair Professorship.
@@ A.1 Extra Nodes
  • $S_{1}$: Nodes with 1-hop and 2-hop neighbours in $G_{i}$.
  • $S_{2}$: Nodes with 1-hop neighbours in $G_{i}$ but $\exists v$ in 2-hop neighbourhood that is not in $G_{i}$.
  • $S_{3}$: Node where $\exists v$ in 1-hop and 2-hop neighbourhood that is not in $G_{i}$.
@@ A.2 Cluster Nodes
Given a partition matrix $P$, the features of the coarsened node are $X^{\prime}=P^{\top}X$. Given a normalized partition matrix, the features of a node $v^{\prime}_{i}\in G^{\prime}$ is the degree-weighted average of the features of nodes in $C_{i}$. This is one of the functions $f$ used to create the features of the cluster node from $C_{i}$.
@@ Appendix B More Algorithms
Algorithm 3 Training GNN (Train on $G^{\prime}$) (Huang et al., 2021)
@@
4:Construct feature matrix for $G^{\prime}$ by $X^{\prime}=P^{\top}X$;
5:Construct labels for $G^{\prime}$ by $Y^{\prime}=\arg\max(P^{\top}Y)$ ;
@@
3:  $X^{(i)}=\sigma(\tilde{D}^{-\frac{1}{2}}\tilde{A}\tilde{D}^{-\frac{1}{2}}X^{(i-1)}\mathcal{W}^{(i-1)})$ $\triangleright$ Equation 1
@@
3:  $X^{(i)}$$\sigma(\tilde{D^{\prime}}^{-\frac{1}{2}}\tilde{A^{\prime}}\tilde{D^{\prime}}^{-\frac{1}{2}}X^{(i-1)}\mathcal{W}^{(i-1)})\triangleright$ Equation 1
@@ C.1 Comparison with Baseline
[…]
@@
    We observe lower inference time complexity of Cluster Node method as compared to Extra Node method under both inference setups. This is expected as $|C_{G_{i}}|\leq|\mathcal{E}_{{G_{i}}}|,\ \forall i\in\{1,2,\dots,k\}$.
C.2 Preprocessing Time Complexity
A comparative analysis of our preprocessing (coarsening) overhead against other state-of-the-art scaling methods has been conducted. As shown in the Table 9 below, our preprocessing time complexity is comparable to, and often faster than, existing SOTA methods. Thus, even when end-to-end latency is considered, FIT-GNN maintains its efficiency advantage.
Table 9: Asymptotic time complexities for different methods. NN is the number of nodes in the original graph, $M$ is the number of edges and $C$ is the number of classes; rest of the symbols follow notation defined in Section 3 and 4. Blue and Red indicate CPU and GPU respectively - the hardware used for computation. Full Graph setup indicates testing on all the test nodes in a graph while Single Node setup represents testing for a single test node.
Method Preprocessing Training Inference (Full Graph) Inference (Single Node) Model-Agnostic
SGGC $M+N$ $k^{2}d+kd^{2}$ $N^{2}d+Nd^{2}$ $N^{2}d+Nd^{2}$
GCOND $\begin{matrix}C(N^{2}+k^{2})d\\ +C(N+k)d^{2}\end{matrix}$ $k^{2}d+kd^{2}$ $N^{2}d+Nd^{2}$ $N^{2}d+Nd^{2}$ $\times$
BONSAI $M+N$ $k^{2}d+kd^{2}$ $N^{2}d+Nd^{2}$ $N^{2}d+Nd^{2}$
FIT-GNN $M+N$ $\begin{matrix}kd^{2}+k^{2}d\\ +\sum_{i=1}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]\end{matrix}$ $\sum_{i=1}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]$ $\max_{i}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]$
We further extend our complexity analysis to the practical scenario, where a new test node $v$ is introduced to the graph $G$. This setting is critical for real-world applications where graphs evolve dynamically. We compare three distinct inference strategies for this new node:
  1. 1.
    Full Graph: Construct $G_{new}$ by adding $v$ to the original full graph $G$ and infer on $G_{new}$ using the pre-trained network.
  2. 2.
    2nd-hop Neighborhood: Sample only the 2nd-hop neighborhood of $v$ from $G$ and infer using the pre-trained network (assuming a 2-layer architecture).
  3. 3.
    FIT-GNN Subgraph: Assign $v$ to a relevant subgraph $G_{i}$ (e.g., the one containing the majority of its 1st-hop neighborhood) and append the necessary Extra/Cluster nodes. Inference is then performed strictly within the modified subgraph $G_{i}$.
We present the time complexity comparison for these methods in the Table 10 below. The most significant advantage of the FIT-GNN approach is that inference complexity depends solely on the subgraph size $\bar{n}_{i}$ ($\bar{n}_{i}\ll n$). In contrast, the Full Graph approach remains tied to the global parameter $n$ due to the necessity of accessing the full adjacency matrix.
Table 10: Time complexity of different inference strategies when a new node $v$ is added to graph $G$
Inference Strategy Preprocessing Complexity Inference Complexity Notes
[…]